# What Differentiates Educational Literature? A Multimodal Fusion Approach of Transformers and Computational Linguistics

### v1

@@ Abstract
The integration of new literature into the English curriculum remains a challenge since educators often lack scalable tools to rapidly evaluate readability and adapt texts for diverse classroom needs. This study proposes to address this gap through a multimodal approach that combines transformer-based text classification with linguistic feature analysis to align texts with UK Key Stages. Eight state-of-the-art Transformers were fine-tuned on segmented text data, with BERT achieving the highest unimodal F1 score of 0.75. In parallel, 500 deep neural network topologies were searched for the classification of linguistic characteristics, achieving an F1 score of 0.392. The fusion of these modalities shows a significant improvement, with every multimodal approach outperforming all unimodal models. In particular, the ELECTRA Transformer fused with the neural network achieved an F1 score of 0.996. The proposed approach is finally encapsulated in a stakeholder-facing web application, providing non-technical stakeholder access to real-time insights on text complexity, reading difficulty, curriculum alignment, and recommendations for learning age range. The application empowers data-driven decision making and reduces manual workload by integrating AI-based recommendations into lesson planning for English literature.
@@ 3.1 Data Collection and Preprocessing
(b) Total books per belonging to a UK Key Stages.

### v3

@@ Abstract
The integration of new literature into the English curriculum remains a challenge since educators often lack scalable tools to rapidly evaluate readability and adapt texts for diverse classroom needs. This study proposes to address this gap through a multimodal approach that combines transformer-based text classification with linguistic feature analysis to align texts with UK Key Stages. Eight state-of-the-art Transformers were fine-tuned on segmented text data, with BERT achieving the highest unimodal F1 score of 0.75. In parallel, 500 deep neural network topologies were searched for the classification of linguistic characteristics, achieving an F1 score of 0.392. The fusion of these modalities shows a significant improvement, with every multimodal approach outperforming all unimodal models. In particular, the ELECTRA Transformer fused with the neural network achieved an F1 score of 0.996. Unimodal and multimodal approaches are shown to have statistically significant differences in all validation metrics (accuracy, precision, recall, F1 score) except for inference time. The proposed approach is finally encapsulated in a stakeholder-facing web application, providing non-technical stakeholder access to real-time insights on text complexity, reading difficulty, curriculum alignment, and recommendations for learning age range. The application empowers data-driven decision making and reduces manual workload by integrating AI-based recommendations into lesson planning for English literature.
@@ 3.1 Data Collection and Preprocessing
(b) Total books belonging to each UK Key Stage.
@@ 3.3 Web Application for Inference and Reporting
3.4 Experimental Setup
The experiments in this work were executed on a server containing 4 Nvidia RTX A2000 GPUs. Each GPU had 3,328 CUDA cores and 12GB of VRAM. For training, all four GPUs were used. For inference, one GPU was used in terms of processing, but it was observed that temporary model storage on the VRAM was shared across each of the four units. All models were trained via the PyTorch library[27], and the educator-facing web application was developed using the Flask framework[28].
@@ 4 Results and Discussion
Table 5: Results of the paired t-test for all metrics when comparing unimodal to multimodal approaches.
Metric t-statistic p-value p <0.05
Accuracy 9.45 <0.001 Yes
Precision 21.95 <0.001 Yes
Recall 9.45 <0.001 Yes
F1 13.30 <0.001 Yes
Inference Time (S) 1.27 0.244 No
Following the observation that all multimodal models seemed to outperform unimodal approaches, further validation is performed through paired t-tests on the performance metrics of accuracy, precision, recall, F1 score and inference time. The results can be observed in Table 5, which shows that these improvements in the classification metrics are statistically significant ($p<0.05$). Furthermore, the paired t-test for inference time ($p=0.244$) shows that there is no statistically significant difference, suggesting that the addition of the artificial neural network does not have a considerable impact on processing time.